# LiDAR Point Cloud-based Multiple Vehicle Tracking with Probabilistic Measurement-Region Association

### v2

@@ II-D Probabilistic Measurement-Region Association
The original DRA method in [14] uses identical prior association probability $P\{\varphi_{k}^{\eta}|\mathbf{Z}^{k-1}\}=1/a_{k}$ for each association, which is not an accurate assumption. A ray-based strategy is proposed by IMM-DRA in [13] to improve the calculation of $P\{\varphi_{k}^{\eta}|\mathbf{Z}^{k-1}\}$. However, since the IMM-DRA is designed for radar, the “visibility” of edge region is not considered in the modeling. Our PMRA model is tailored to LiDAR sensors, which utilizes the visible angle of different extent regions to calculate the prior association probability $P\{\varphi_{k}^{\eta}|\mathbf{Z}^{k-1}\}$ more accurately. We further specify the $\eta$-th measurement-region association at time $k$ as $\varphi_{k}^{\eta}=\{\phi_{k}^{\eta,m}\}_{m=1}^{|\mathbf{Y}_{k}|}$, where $\phi_{k}^{\eta,m}$ is the region assigned for the $m$-th measurement. By assuming that the measurements in $\mathbf{Y}_{k}$ are mutually independent, we have
@@
$P_{1}=\frac{P_{\mathrm{vis}}b_{\phi_{k}^{\eta,m}}(\mathbf{x}_{k|k-1})\theta_{\phi_{k}^{\eta,m}}(\mathbf{x}_{k|k-1})}{\sum_{{n^{\prime}}=1}^{4}b_{r_{n^{\prime}}}(\mathbf{x}_{x|k-1})\theta_{r_{n^{\prime}}}(\mathbf{x}_{k|k-1})}$ (22b)
$P_{2}=\frac{P_{\mathrm{inv}}(1-b_{\phi_{k}^{\eta,m}}(\mathbf{x}_{k|k-1}))\theta_{\phi_{k}^{\eta,m}}(\mathbf{x}_{k|k-1})}{\sum_{{n^{\prime}}=1}^{4}(1-b_{r_{n^{\prime}}}(\mathbf{x}_{x|k-1}))\theta_{r_{n^{\prime}}}(\mathbf{x}_{k|k-1})}$ (22c)
@@ III Particle-based Implementation of the Proposed Method
For the system model in Section II, a closed-form filtering recursion of the posterior PMBM density is difficult to obtain. Moreover, the standard linearization methods such as Taylor expansion and UT may not yield a feasible estimator due to the highly nonlinear eigen-decomposition required in the measurement model. Therefore, a particle-based implementation of the proposed PMRA-PMBM algorithm is considered in this section. The distribution of kinematic state $\mathbf{m}_{k}$ and extent state $E_{k}$ of a target is represented by a weighted set of particles $\Xi_{k}=\{(\mathbf{m}_{k}^{(l)},E_{k}^{(l)},w_{k}^{(l)})\}_{l=1}^{L}$, where $\sum_{l}w_{k}^{(l)}=1$. Note that a closed-form recursion based on gamma conjugate prior is used to estimate the measurement rate and $\gamma_{k}$ is not contained in the particles. Consequently, the distribution of state $\mathbf{x}_{k}$ can be represented by the PDF $\mathcal{F}(\mathbf{x}_{k};\alpha_{k},\beta_{k},\Xi_{k})$.
@@ III-A Data Association Problem
Since the actual origins of measurements are unknown, the data association problem must be handled in the PMRA-PMBM algorithm. At time $k$, let the global hypotheses indexed by set $\mathbf{J}_{k}$ and denote the index set of existing targets in the $j$-th global hypothesis by $\mathbf{I}^{j}_{k}$. The measurement index set satisfies $\mathbf{M}_{k}\cap\mathbf{I}^{j}_{k}=\emptyset$ for all $j\in\mathbf{J}_{k}$. Let $\mathbb{A}^{j}_{k}$ be the space of all data associations in the $j$-th global hypothesis. Then, a data association $\mathbf{A}\in\mathbb{A}^{j}_{k}$ can be defined as partitioning the union set $\mathbf{M}_{k}\cup\mathbf{I}^{j}_{k}$ into disjoint nonempty subsets $\{\mathbf{S}\}$ that satisfy $\uplus_{\mathbf{S}\in\mathbf{A}}\mathbf{S}=\mathbf{M}_{k}\cup\mathbf{I}^{j}_{k}$ [11]. A subset $\mathbf{S}$ is called an index cell, which can contain any measurement indices $m\in\mathbf{M}_{k}$ and at most one target index $i_{\mathbf{S}}\in\mathbf{I}^{j}_{k}$, because a measurement cannot originate from more than one target in our system assumption. The measurements in index cell $\mathbf{S}$ are denoted by $\mathbf{Y}_{\mathbf{S}}=\{\mathbf{z}^{m}_{k}\}_{m\in\mathbf{S}\cap\mathbf{M}_{k}}$. Thus, if $i_{\mathbf{S}}\in\mathbf{S}$ and $\mathbf{Y}_{\mathbf{S}}\neq\emptyset$, $\mathbf{S}$ assigns the measurements $\mathbf{Y}_{\mathbf{S}}$ to the target $i_{\mathbf{S}}$; if $\mathbf{S}$ contains no target index, it assigns $\mathbf{Y}_{\mathbf{S}}$ to a newborn target or clutter; if $\mathbf{S}=\{i_{\mathbf{S}}\}$, it means that target $i_{\mathbf{S}}$ is not detected.
The computational complexity of the data association problem can be significantly reduced by eliminating improbable association hypotheses with clustering and gating. In the PMRA-PMBM filter, the measurements $\mathbf{Z}_{k}$ are first partitioned into disjoint clusters by the DBSCAN algorithm [24]. Next, a gating step is performed to identify the associations between clusters and targets. If a cluster $\mathbf{Y}\subseteq\mathbf{Z}_{k}$ contains measurements within a certain distance $d_{\mathrm{in}}$ around the predicted position of the $i$-th existing target, an index cell $\mathbf{S}$ can be formed by assigning $\mathbf{Y}_{\mathbf{S}}=\mathbf{Y}$ and $i_{\mathbf{S}}=i$. If the minimum distance from the center of a cluster $\mathbf{Y}^{b}\subseteq\mathbf{Z}_{k}$ to the predicted position of any existing target exceeds a certain threshold $d_{\mathrm{out}}$, the cluster is possibly originating from a newborn target. According to the clustering and gating results denoted by $\mathcal{CG}_{k}$, only a subset of all possible data associations is calculated at each time step. For further discussions on data association, clustering, and gating see, e.g., [9, 11, 4].
@@ III-B Target State Initialization
The state of a newborn target is initialized when a new Poisson component is added to the PMBM density to represent a birth event. Each Poisson component models the state of an undetected target, and the PPP intensity $D^{u}(\mathbf{x})$ is maintained as the weighted sum of Poisson components. Specifically, $D^{u}(\mathbf{x})=\sum_{{h\in\mathbf{H}^{u}}}w^{u,h}\mathcal{F}(\mathbf{x};\alpha^{u,h},\beta^{u,h},\Xi^{u,h})$, where $w^{u,h}>0$ is the component weight, and $\mathbf{H}^{u}$ is an index set.
At time $k$, assume a set of measurements $\mathbf{Y}_{k}^{b}$ is used to initialize the state of a newborn target $\{\alpha^{b}_{k},\beta^{b}_{k},\Xi^{b}_{k}\}$. Then, the gamma parameters of the measurement rate $\gamma_{k}$ are assigned with predefined values $\alpha^{b}_{k}=\alpha^{b}$, $\beta^{b}_{k}=\beta^{b}$, and the kinematic part of particles $\Xi^{b}_{k}$ is obtained by sampling from the following proposal distributions
$\mathbf{m}_{k}^{b,(l)}=\left[x_{k}^{b,(l)},y_{k}^{b,(l)},\dot{x}_{k}^{b,(l)},\dot{y}_{k}^{b,(l)},\omega_{k}^{b,(l)}\right]^{\mathrm{T}}$ (23a)
$[x_{k}^{b,(l)},y_{k}^{b,(l)}]^{\mathrm{T}}\sim\mathcal{N}(\bar{\mathbf{z}}_{k},Q_{\mathrm{pos}})$ (23b)
$[\dot{x}_{k}^{b,(l)},\dot{y}_{k}^{b,(l)}]^{\mathrm{T}}\sim\mathcal{N}(\dot{\mathbf{m}}_{k},Q_{\mathrm{vel}})$ (23c)
$\omega_{k}^{b,(l)}\sim\mathcal{N}(0,Q_{\mathrm{turn}})$ (23d)
where $\bar{\mathbf{z}}_{k}=\sum_{\mathbf{z}\in\mathbf{Y}_{k}^{b}}\mathbf{z}/|\mathbf{Y}_{k}^{b}|$ is the approximated target position, $\dot{\mathbf{m}}_{k}$ is the approximated target velocity, $Q_{\mathrm{pos}}$, $Q_{\mathrm{vel}}$, and $Q_{\mathrm{turn}}$ are predefined covariance matrices.
The extent part of $\Xi^{b}_{k}$ is sampled from an inverse Wishart distribution with DOF parameter $q^{b}$ and scale matrix $V^{b}_{k}$:
$E_{k}^{b,(l)}\sim\mathcal{IW}(q^{b},V^{b}_{k})$ (24a)
$V^{b}_{k}=(q^{b}-3)\mathrm{Rot}(\dot{\mathbf{m}}_{k})\begin{bmatrix}\bar{e}_{1}&0\\ 0&\bar{e}_{2}\end{bmatrix}\mathrm{Rot}(\dot{\mathbf{m}}_{k})^{\mathrm{T}}$ (24d)
@@ III-C PMRA-PMBM Prediction
In the particle-based implementation of PMRA-PMBM, the state transition PDF defined in (11) is represented by $\mathcal{F}(\mathbf{x}_{+};\alpha_{+},\beta_{+},\Xi_{+})$ where the subscript $+$ is the abbreviation of $k|k-1$. The predicted gamma parameters $\alpha_{+}$, $\beta_{+}$ are given in (II-B), and the predicted particles and weights are
@@
$D_{+}^{u},\{(C_{+}^{j},\{(r_{+}^{j,i},f_{+}^{j,i})\}_{i\in\mathbf{I}_{k-1}^{j}})\}_{j\in\mathbf{J}_{k-1}}.$ (27)
Based on the assumptions in Table I, the pseudo code of PMRA-PMBM prediction is provided in Table II, where the time index $k-1$ is omitted for notational simplicity. The predicted PMBM density is obtained by applying the state transition model (II-B) and (III-C) to PPP and MBM densities. Different from the standard PMBM implementations (e.g.,[11]), the predicted PPP intensity $D_{+}^{u}$ of PMRA-PMBM does not include a predefined birth intensity $D^{b}_{+}$ with known parameters. Instead, the intensity of newborn target is calculated in the update procedure with clustering and gating results.
@@
1. Empty initial PPP: $\mathbf{H}^{u}_{0}=\emptyset$.
@@
4. Clutter Poisson rate $\mu^{c}$ is known and the spatial distribution is $f^{c}(\mathbf{z})=1/\mathrm{V}$, where $\mathrm{V}$ is the volume of the surveillance region.
@@
Input: $D^{u},\{(C^{j},\{(r^{j,i},f^{j,i})\}_{i\in\mathbf{I}^{j}})\}_{j\in\mathbf{J}}$.
Output: $D_{+}^{u},\{(C_{+}^{j},\{(r_{+}^{j,i},f_{+}^{j,i})\}_{i\in\mathbf{I}^{j}})\}_{j\in\mathbf{J}}$.
Initialize: $D_{+}^{u}\leftarrow 0$
for $h\in\mathbf{H}^{u}$ do:
 From $\alpha^{u,h}$, $\beta^{u,h}$, $\Xi^{u,h}$, compute $\alpha_{+}^{u,h}$, $\beta_{+}^{u,h}$, $\Xi_{+}^{u,h}$ by (II-B), (III-C).
 Increment: $D^{u}_{+}\leftarrow D^{u}_{+}+w^{u,m}\mathcal{F}(\alpha_{+}^{u,h},\beta_{+}^{u,h},\Xi_{+}^{u,h})$
@@
  From $\alpha^{j,i}$, $\beta^{j,i}$, $\Xi^{j,i}$, compute $\alpha_{+}^{j,i}$, $\beta_{+}^{j,i}$, $\Xi_{+}^{j,i}$ by (II-B), (III-C).

### v3

@@ II-D Probabilistic Measurement-Region Association
The original DRA method in [14] uses identical prior association probability $P\{\varphi_{k}^{\eta}|\mathbf{Z}^{k-1}\}=1/a_{k}$ for each association, which is not an accurate assumption. A ray-based strategy is proposed by IMM-DRA in [13] to improve the calculation of $P\{\varphi_{k}^{\eta}|\mathbf{Z}^{k-1}\}$. However, since the IMM-DRA is designed for radar, the “visibility” of edge regions is not considered in the modeling. Our PMRA model is tailored to LiDAR sensors, which utilizes the visible angle of different extent regions to calculate the prior association probability $P\{\varphi_{k}^{\eta}|\mathbf{Z}^{k-1}\}$ more accurately. We further specify the $\eta$-th measurement-region association at time $k$ as $\varphi_{k}^{\eta}=\{\phi_{k}^{\eta,m}\}_{m=1}^{|\mathbf{Y}_{k}|}$, where $\phi_{k}^{\eta,m}$ is the region assigned for the $m$-th measurement. By assuming that the measurements in $\mathbf{Y}_{k}$ are mutually independent, we have
@@
$P_{1}=\frac{P_{\mathrm{vis}}b_{\phi_{k}^{\eta,m}}(\mathbf{x}_{k|k-1})\theta_{\phi_{k}^{\eta,m}}(\mathbf{x}_{k|k-1})}{\sum_{{n^{\prime}}=1}^{4}b_{r_{n^{\prime}}}(\mathbf{x}_{k|k-1})\theta_{r_{n^{\prime}}}(\mathbf{x}_{k|k-1})}$ (22b)
$P_{2}=\frac{P_{\mathrm{inv}}(1-b_{\phi_{k}^{\eta,m}}(\mathbf{x}_{k|k-1}))\theta_{\phi_{k}^{\eta,m}}(\mathbf{x}_{k|k-1})}{\sum_{{n^{\prime}}=1}^{4}(1-b_{r_{n^{\prime}}}(\mathbf{x}_{k|k-1}))\theta_{r_{n^{\prime}}}(\mathbf{x}_{k|k-1})}$ (22c)
@@ III Particle-based Implementation of the Proposed Method
For the system model in Section II, a closed-form filtering recursion of the posterior PMBM density is difficult to obtain. Moreover, the standard linearization methods such as Taylor expansion and UT may not yield a feasible estimator due to the highly nonlinear eigen-decomposition required in the measurement model. Therefore, a particle-based implementation of the proposed PMRA-PMBM algorithm is considered in this section. The distribution of kinematic state $\mathbf{m}_{k}$ and extent state $E_{k}$ of a target is represented by a weighted set of particles $\Xi_{k}=\{(\mathbf{m}_{k}^{(l)},E_{k}^{(l)},w_{k}^{(l)})\}_{l=1}^{L}$, where $\sum_{l}w_{k}^{(l)}=1$. Note that a closed-form recursion based on the gamma conjugate prior is used to estimate the measurement rate, and $\gamma_{k}$ is not contained in the particles. Consequently, the distribution of state $\mathbf{x}_{k}$ can be represented by the PDF $\mathcal{F}(\mathbf{x}_{k};\alpha_{k},\beta_{k},\Xi_{k})$.
@@ III-A Data Association Problem
Since the actual origins of measurements are unknown, the data association problem must be handled in the PMRA-PMBM algorithm. At time $k$, let the global hypotheses be indexed by set $\mathbf{J}_{k}$ and denote the index set of existing targets in the $j$-th global hypothesis by $\mathbf{I}^{j}_{k}$. The measurement index set satisfies $\mathbf{M}_{k}\cap\mathbf{I}^{j}_{k}=\emptyset$ for all $j\in\mathbf{J}_{k}$. Let $\mathbb{A}^{j}_{k}$ be the space of all data associations in the $j$-th global hypothesis. Then, a data association $\mathbf{A}\in\mathbb{A}^{j}_{k}$ can be defined as partitioning the union set $\mathbf{M}_{k}\cup\mathbf{I}^{j}_{k}$ into disjoint nonempty subsets $\{\mathbf{S}\}$ that satisfy $\uplus_{\mathbf{S}\in\mathbf{A}}\mathbf{S}=\mathbf{M}_{k}\cup\mathbf{I}^{j}_{k}$ [11]. A subset $\mathbf{S}$ is called an index cell, which can contain any measurement indices $m\in\mathbf{M}_{k}$ and at most one target index $i_{\mathbf{S}}\in\mathbf{I}^{j}_{k}$ because a measurement cannot originate from more than one target in our system assumption. The measurements in index cell $\mathbf{S}$ are denoted by $\mathbf{Y}_{\mathbf{S}}=\{\mathbf{z}^{m}_{k}\}_{m\in\mathbf{S}\cap\mathbf{M}_{k}}$. Thus, if $i_{\mathbf{S}}\in\mathbf{S}$ and $\mathbf{Y}_{\mathbf{S}}\neq\emptyset$, $\mathbf{S}$ assigns the measurements $\mathbf{Y}_{\mathbf{S}}$ to the target $i_{\mathbf{S}}$; if $\mathbf{S}$ contains no target index, it assigns $\mathbf{Y}_{\mathbf{S}}$ to a newborn target or clutters; if $\mathbf{S}=\{i_{\mathbf{S}}\}$, it means that the target $i_{\mathbf{S}}$ is not detected.
The computational complexity of the data association can be significantly reduced by eliminating improbable association hypotheses with clustering and gating. In the PMRA-PMBM filter, the measurements $\mathbf{Z}_{k}$ are first partitioned into disjoint clusters by the DBSCAN algorithm [24]. Next, a gating step is performed to identify the associations between clusters and targets. If a cluster $\mathbf{Y}\subseteq\mathbf{Z}_{k}$ contains measurements within a certain distance $d_{\mathrm{in}}$ around the predicted position of the $i$-th existing target, an index cell $\mathbf{S}$ can be formed by assigning $\mathbf{Y}_{\mathbf{S}}=\mathbf{Y}$ and $i_{\mathbf{S}}=i$. If the minimum distance from the center of a cluster $\mathbf{Y}^{\mathrm{b}}\subseteq\mathbf{Z}_{k}$ to the predicted position of any existing target exceeds a certain threshold $d_{\mathrm{out}}$, the cluster is possibly originating from a newborn target or clutters. According to the clustering and gating results denoted by $\mathcal{CG}_{k}$, only a subset of all possible data associations is calculated at each time step. For further discussions on data association, clustering, and gating, see, e.g., [4, 9], and [11].
@@ III-B Target State Initialization
The state of a newborn target is initialized when a new Poisson component is added to the PMBM density to represent a birth event. Each Poisson component models the state of an undetected target, and the PPP intensity $D^{\mathrm{u}}(\mathbf{x})$ is maintained as the weighted sum of Poisson components. Specifically, $D^{\mathrm{u}}(\mathbf{x})=\sum_{{h\in\mathbf{H}^{\mathrm{u}}}}w^{\mathrm{u},h}\mathcal{F}(\mathbf{x};\alpha^{\mathrm{u},h},\beta^{\mathrm{u},h},\Xi^{\mathrm{u},h})$, where $w^{\mathrm{u},h}>0$ is the component weight, and $\mathbf{H}^{\mathrm{u}}$ is an index set.
At time $k$, assume a set of measurements $\mathbf{Y}_{k}^{\mathrm{b}}$ is used to initialize the state of a newborn target $\{\alpha^{\mathrm{b}}_{k},\beta^{\mathrm{b}}_{k},\Xi^{\mathrm{b}}_{k}\}$. Then, the gamma parameters of the measurement rate $\gamma_{k}$ are assigned with predefined values $\alpha^{\mathrm{b}}_{k}=\alpha^{\mathrm{b}}$, $\beta^{\mathrm{b}}_{k}=\beta^{\mathrm{b}}$, and the kinematic part of particles $\Xi^{\mathrm{b}}_{k}$ is obtained by sampling from the following proposal distributions
$\mathbf{m}_{k}^{\mathrm{b},(l)}=\left[x_{k}^{\mathrm{b},(l)},\dot{x}_{k}^{\mathrm{b},(l)},y_{k}^{\mathrm{b},(l)},\dot{y}_{k}^{\mathrm{b},(l)},\omega_{k}^{\mathrm{b},(l)}\right]^{\mathrm{T}}$ (23a)
$[x_{k}^{\mathrm{b},(l)},y_{k}^{\mathrm{b},(l)}]^{\mathrm{T}}\sim\mathcal{N}(\bar{\mathbf{z}}_{k},Q_{\mathrm{pos}})$ (23b)
$[\dot{x}_{k}^{\mathrm{b},(l)},\dot{y}_{k}^{\mathrm{b},(l)}]^{\mathrm{T}}\sim\mathcal{N}(\dot{\mathbf{m}}_{k},Q_{\mathrm{vel}})$ (23c)
$\omega_{k}^{\mathrm{b},(l)}\sim\mathcal{N}(0,Q_{\mathrm{turn}})$ (23d)
where $\bar{\mathbf{z}}_{k}=\sum_{\mathbf{z}\in\mathbf{Y}_{k}^{\mathrm{b}}}\mathbf{z}/|\mathbf{Y}_{k}^{\mathrm{b}}|$ is the approximated target position, $\dot{\mathbf{m}}_{k}$ is the approximated target velocity, $Q_{\mathrm{pos}}$, $Q_{\mathrm{vel}}$, and $Q_{\mathrm{turn}}$ are predefined covariance matrices.
The extent part of $\Xi^{\mathrm{b}}_{k}$ is sampled from an inverse Wishart distribution with DOF parameter $q^{\mathrm{b}}$ and scale matrix $V^{\mathrm{b}}_{k}$:
$E_{k}^{\mathrm{b},(l)}\sim\mathcal{IW}(q^{\mathrm{b}},V^{\mathrm{b}}_{k})$ (24a)
$V^{\mathrm{b}}_{k}=(q^{\mathrm{b}}-3)\mathrm{Rot}(\dot{\mathbf{m}}_{k})\begin{bmatrix}\bar{e}_{1}&0\\ 0&\bar{e}_{2}\end{bmatrix}\mathrm{Rot}(\dot{\mathbf{m}}_{k})^{\mathrm{T}}$ (24b)
@@ III-C PMRA-PMBM Prediction
In the particle-based implementation of PMRA-PMBM, the state transition PDF defined in (11) is represented by $\mathcal{F}(\mathbf{x}_{+};\alpha_{+},\beta_{+},\Xi_{+})$ where the subscript $+$ is the abbreviation of $k|k-1$. The predicted gamma parameters $\alpha_{+}$, $\beta_{+}$ are given in (10), and the predicted particles and weights are
@@
$D_{+}^{\mathrm{u}},\{(C_{+}^{j},\{(r_{+}^{j,i},f_{+}^{j,i})\}_{i\in\mathbf{I}_{k-1}^{j}})\}_{j\in\mathbf{J}_{k-1}}.$ (27)
Based on the assumptions in Table I, the pseudo code of PMRA-PMBM prediction is provided in Table II, where the time index $k-1$ is omitted for notational simplicity. The predicted PMBM density is obtained by applying the state transition model (10) and (26) to the PPP and MBM densities. Different from the standard PMBM implementations (e.g.,[11]), the predicted PPP intensity $D_{+}^{\mathrm{u}}$ of PMRA-PMBM does not include a predefined birth intensity $D^{\mathrm{b}}_{+}$ with known parameters. Instead, the intensity of newborn target is calculated in the update procedure with clustering and gating results.
@@
1. Empty initial PPP: $\mathbf{H}^{\mathrm{u}}_{0}=\emptyset$.
@@
4. Clutter Poisson rate $\mu^{\mathrm{c}}$ is known and the spatial distribution is $f^{\mathrm{c}}(\mathbf{z})=1/\mathrm{V}$, where $\mathrm{V}$ is the volume of the surveillance region.
@@
Input: $D^{\mathrm{u}},\{(C^{j},\{(r^{j,i},f^{j,i})\}_{i\in\mathbf{I}^{j}})\}_{j\in\mathbf{J}}$.
Output: $D_{+}^{\mathrm{u}},\{(C_{+}^{j},\{(r_{+}^{j,i},f_{+}^{j,i})\}_{i\in\mathbf{I}^{j}})\}_{j\in\mathbf{J}}$.
Initialize: $D_{+}^{\mathrm{u}}\leftarrow 0$.
for $h\in\mathbf{H}^{\mathrm{u}}$ do:
 From $\alpha^{\mathrm{u},h}$, $\beta^{\mathrm{u},h}$, $\Xi^{\mathrm{u},h}$, compute $\alpha_{+}^{\mathrm{u},h}$, $\beta_{+}^{\mathrm{u},h}$, $\Xi_{+}^{\mathrm{u},h}$ by (10), (26).
 Increment: $D^{\mathrm{u}}_{+}\leftarrow D^{\mathrm{u}}_{+}+w^{\mathrm{u},m}\mathcal{F}(\alpha_{+}^{\mathrm{u},h},\beta_{+}^{\mathrm{u},h},\Xi_{+}^{\mathrm{u},h})$.
@@
  From $\alpha^{j,i}$, $\beta^{j,i}$, $\Xi^{j,i}$, compute $\alpha_{+}^{j,i}$, $\beta_{+}^{j,i}$, $\Xi_{+}^{j,i}$ by (10), (26).